OpenStudy (cenaida):

The sum of five times a number and twice a second number is 18, but the difference of three times the first number and seven times the second number is -22. Find the two numbers using the elimination method. A. -2 and -4 B. -2 and 4 C. 2 and -4 D. 2 and 4

OpenStudy (misty1212):

the sum of five times a number lets call it \(x\) so the sum of \(5x\) and two times a second number, lets call it \(y\) so \[5x+2y\] is \(18\)or \[5x+2y=18\]

OpenStudy (misty1212):

the difference of 3 times the first number and seven times the second number is \(-22\) translates in to \[3x-7y=-22\]

OpenStudy (misty1212):

your job it to solve \[5x+2y=18\\ 3x-7y=-22\]

OpenStudy (misty1212):

do you know how to do that?

OpenStudy (cenaida):

no

OpenStudy (shamim):

Just let one number is =x Another number is=y

OpenStudy (misty1212):

lol then you can't solve this problem you can do it by hand or by machine

OpenStudy (misty1212):

you pick

OpenStudy (cenaida):

is it D?

OpenStudy (anonymous):

\[\{5 a+2 b=18, 3 a-7 b=-22\} , \{a\to 2, b\to 4\} \]

OpenStudy (shamim):

Fivs times of a number x=5x Right?